OpenStudy (anonymous):

(square root of 5+4)(square root of 5-1)

OpenStudy (anonymous):

\[\sqrt{5+4}\sqrt{5-1}\]

OpenStudy (diyadiya):

\[\sqrt{5+4}=\sqrt{9}\] Now what is \[\sqrt{5-1} \ ?\]

OpenStudy (hba):

This is difficult.

OpenStudy (anonymous):

\[\sqrt{4}\]

OpenStudy (diyadiya):

@iknott1 Right! \[\sqrt{9}=3\]\[\sqrt{4}=2\] so \[3 \times 2 = \ ?\]

OpenStudy (anonymous):

6

OpenStudy (diyadiya):

RIght! :)